# Efficient Streaming Language Models with Attention Sinks

### 12 Dec 2023 v3

@@ 1 Introduction
Finally, we confirm our attention sink hypothesis and demonstrate that language models can be pre-trained to require only a single attention sink token for streaming deployment. Specifically, we suggest that an extra learnable token at the beginning of all training samples can serve as a designated attention sink. By pre-training 160-million parameter language models from scratch, we demonstrate that adding this single sink token preserves the model’s performance in streaming cases. This stands in contrast to vanilla models, which necessitate the reintroduction of multiple initial tokens as attention sinks to achieve the same performance level.
@@ 3.1 The Failure of Window Attention and Attention Sinks
Figure 3 shows the perplexity of language modeling on a 20K token text. It is evident that perplexity spikes when the text length surpasses the cache size, led by the exclusion of initial tokens. This suggests that the initial tokens, regardless of their distance from the tokens being predicted, are crucial for maintaining the stability of LLMs.
@@ 3.3 Pre-Training LLMs with Attention Sinks
Table 3: Comparison of vanilla attention with prepending a zero token and a learnable sink token during pre-training. To ensure stable streaming perplexity, the vanilla model required several initial tokens. While Zero Sink demonstrated a slight improvement, it still needed other initial tokens. Conversely, the model trained with a learnable Sink Token showed stable streaming perplexity with only the sink token added. Cache config $x$+$y$ denotes adding $x$ initial tokens with $y$ recent tokens. Perplexity is evaluated on the first sample in the PG19 test set.
@@
As elaborated in Section 3.1, a significant reason for the model’s excessive attention to multiple initial tokens is the absence of a designated sink token to offload excessive attention scores. Due to this, the model inadvertently designates globally visible tokens, primarily the initial ones, as attention sinks. A potential remedy can be the intentional inclusion of a global trainable attention sink token, denoted as a “Sink Token", which would serve as a repository for unnecessary attention scores. Alternatively, replacing the conventional SoftMax function with a variant like SoftMax-off-by-One (Miller, 2023),
@@
which does not require the attention scores on all contextual tokens to sum up to one, might also be effective. Note that this SoftMax alternative is equivalent to using a token with an all-zero Key and Value features in the attention computation. We denote this method as “Zero Sink" to fit it consistently in our framework.
@@ 4.1 Language Modeling on Long Texts Across LLM Families and Scales
Figure 5: Language modeling perplexity of StreamingLLM for super long texts with 4 million tokens across various LLM families and model scales. The perplexity remains stable throughout. We use the concatenated test set of PG19 (100 books) to perform language modeling, with perplexity fluctuations attributable to the transition between books.
@@ 4.4 Ablation Studies
Table 6: Effects of cache size on StreamingLLM’s performance. Increasing the cache size in StreamingLLM doesn’t consistently yield a decrease in perplexity across all models, suggesting these models may not be fully exploiting the provided context. Cache config $x$+$y$ denotes adding $x$ initial tokens with $y$ recent tokens. Perplexity is evaluated on 400K tokens in the concatenated PG19 test set.
@@ 4.5 Efficency Results
We benchmark its decoding latency and memory usage against the sliding window with re-computation, which is the only baseline with acceptable performance. Both methods are implemented using the Huggingface Transformers library (Wolf et al., 2020) and tested on a single NVIDIA A6000 GPU using the Llama-2-7B and Llama-2-13B models. As depicted in Figure 10, as the cache size increases, StreamingLLM’s decoding speed demonstrates a linear growth. The sliding window with re-computation baseline has a quadratic rise in decoding latency. Thus, StreamingLLM achieves an impressive speedup, reaching up to 22.2$\times$ per token. Despite its reduced latency, StreamingLLM sustains a memory footprint consistent with the re-computation baseline.

### 07 Apr 2024 v4

@@ 1 Introduction
Furthermore, we confirm our attention sink hypothesis and demonstrate that language models can be pre-trained to require only a single attention sink token for streaming deployment. Specifically, we suggest that an extra learnable token at the beginning of all training samples can serve as a designated attention sink. By pre-training 160-million parameter language models from scratch, we demonstrate that adding this single sink token preserves the model’s performance in streaming cases. This stands in contrast to vanilla models, which necessitate the reintroduction of multiple initial tokens as attention sinks to achieve the same performance level.
@@
Finally, we emphasize that StreamingLLM efficiently generates coherent text from tokens within the KV cache without extending the LLMs’ context length. It suits continuous operation needs with minimal memory use and past data reliance. Additionally, StreamingLLM can complement context extension methods to increase the attendable recent context.
@@ 3.1 The Failure of Window Attention and Attention Sinks
Figure 3 shows the perplexity of language modeling on a 20K token text. It is evident that perplexity spikes when the text length surpasses the cache size, led by the exclusion of initial tokens. This suggests that the initial tokens, regardless of their distance from the predicted tokens, are crucial for maintaining the stability of LLMs.
@@ 3.3 Pre-Training LLMs with Attention Sinks
Table 3: Comparison of vanilla attention with prepending a zero token and a learnable sink token during pre-training. To ensure stable streaming perplexity, the vanilla model requires several initial tokens. While Zero Sink shows a slight improvement, it still needs other initial tokens. Conversely, the model trained with a learnable Sink Token shows stable streaming perplexity with only the sink token added. Cache config $x$+$y$ denotes adding $x$ initial tokens with $y$ recent tokens. Perplexity is evaluated on the first sample in the PG19 test set.
@@
As elaborated in Section 3.1, a significant reason for the model’s excessive attention to multiple initial tokens is the absence of a designated sink token to offload excessive attention scores. Due to this, the model inadvertently uses globally visible tokens, primarily the initial ones, as attention sinks. A potential remedy can be the intentional inclusion of a global trainable attention sink token, denoted as a “Sink Token”, which would serve as a repository for unnecessary attention scores. Alternatively, replacing the conventional SoftMax function with a variant like SoftMax-off-by-One (Miller, 2023),
@@
which does not require the attention scores on all contextual tokens to sum up to one, may also be effective. Note that $\text{SoftMax}_{1}$ is equivalent to prepending a token with an all-zero Key and Value features in the attention computation. We denote this method as “Zero Sink” to fit our framework.
@@ 4.1 Language Modeling on Long Texts Across LLM Families and Scales
Figure 5: Language modeling perplexity of StreamingLLM on super long texts with 4 million tokens across various LLM families and scales. The perplexity remains stable throughout. We use the concatenated test set of PG19 (100 books) to perform language modeling, with perplexity fluctuations due to book transitions.
@@ 4.4 Ablation Studies
Table 6: Effects of cache size on StreamingLLM’s performance. Increasing the cache size in StreamingLLM doesn’t consistently yield a decrease in perplexity, showing these models may not fully utilize the provided context. Cache config $x$+$y$ denotes adding $x$ initial tokens with $y$ recent tokens. Perplexity is evaluated on 400K tokens in the concatenated PG19 test set.
@@ 4.5 Efficency Results
We benchmark StreamingLLM’s decoding latency and memory usage against the sliding window with re-computation, which is the only baseline with acceptable quality. Both methods are implemented using the Huggingface Transformers library (Wolf et al., 2020) and tested on a single NVIDIA A6000 GPU using the Llama-2-7B and Llama-2-13B models. As shown in Figure 10, as the cache size increases, StreamingLLM’s decoding speed has a linear growth. The sliding window with re-computation baseline has a quadratic rise in decoding latency. Thus, StreamingLLM achieves an impressive speedup, reaching up to 22.2$\times$ per token. Despite its reduced latency, StreamingLLM sustains a memory footprint consistent with the re-computation baseline.
@@ 5 Conclusion
Reproducibility Statement
All findings presented in this paper can be reproduced. We have made our code and datasets available in this github repo. The models used in this paper are all openly available, and we provide references to access them. Details regarding our experiments, including hyperparameters, training protocols, and evaluation methods, can be found in the Experiments section (Section 4). We are confident that with the provided resources, readers can reproduce the entirety of our presented results.
Impact Statement
StreamingLLM has been widely adopted by various LLM serving solutions including NVIDIA TensorRT-LLM, Intel Extension for Transformers, HuggingFace Transformers, MLC LLM, etc.